Suppose a firm finds that the marginal product of capital is 60 and the marginal product of labor is 20. If the price of capital is $6 and the price of labor is $2.50, describe how the firm should adjust its mix of capital and labor? What will be the result?


1
Expert's answer
2021-07-14T13:24:17-0400

marginal product of capital(MPK)=60

marginal product of labor (MPL)=20

wage rate(w)=$2.50

price of capital(r)=$6

A firm employing labor and capital as input the best combination occur at


"\frac{MPL}{W}=\frac{MPK}{r}"


"\frac{MPL}{W}=\frac{20}{2.50}\\\\\\"


"\frac{MPL}{W}=8"


"\frac{MPK}{r}=\frac{60}{6
}"


"\frac{MPK}{r}=10"


now


"\frac{MPL}{W} <\frac{MPK}{r}\\"


"8<10"

For a firm to attain the best input mix,it should hire more labor and hire less capital so that marginal product of labor increase and marginal product of capital reduces until the equilibrium

is attained.